# A compact Optical Liquid Argon Facility at Roma Tre

### v1

@@ 1 Introduction
Liquid argon (LAr) has been used in neutrino and dark matter experiments as an active medium because of its excellent properties in charge yield and transport, as well as its capacity as a scintillator [1]. To detect the LAr scintillation photons of the vacuum ultraviolet (VUV) wavelength range, standard method uses wavelength shifting materials to produce secondary photons in the visible wavelength range to which typical photomultipliers have the sensitivity.
@@
At Roma Tre university, we are preparing the Optical Liquid Argon Facility (OLAF), with a cylindrical volume of 40 L LAr to host VUV SiPMs and photon sources inside. The first of the two main goals of the facility is to characterize the optical properties of LAr as a scintillation detector to be used in LEGEND-1000 experiment [6] as active shielding. Secondly, the facility is expected to function as a test bench for the R&D of LAr scintillation detector in general. The compact size of OLAF is important for a fast turnaround time to test different configurations and designs of the measurement. In the next section we will introduce the experimental setup in detail.
@@ 2.1 Cryogenics for LAr
The central part of the cryogenic system is a cylindrical vessel made in stainless steel, about 2 m high as shown in figure 2. The innermost vessel can host LAr up to 40 L, surrounded by a jacketed volume of $\sim$20 L where the LN2 is filled. An external vacuum layer then thermally isolates the inner volumes from the surroundings. The three layers are welded together at the top part of the vessels, as shown in the photo in figure 2a. A 200 L Dewar is used to continuously provide LN2 to OLAF.
@@
The thermally insulated pipelines were equipped, as shown in figure 2b, to vent the cold gas to the roof outside the laboratory. The top of the LAr volume is fitted with a transparent sight glass through which the interior can be visually inspected. Temperature sensors placed at multiple vertical locations indicate the liquid level of the LAr and LN2 volumes. The feed-through flanges as shown in figure 2a provide separate sockets for the SiPMs and the temperature sensors.
@@
When the LN2 volume of OLAF is filled, the thermal contact of the jacketed wall provides a cooling power of a few kilowatts to the argon volume. By controlling the inflow of the argon gas, we can produce tens of liters of LAr in a few hours  [^3] This includes the time required to cool the vessels from room temperature with LN2. At 1 atm, the boiling temperature of LN2 (77 K) is lower than the freezing temperature of LAr (83 K). To prevent LAr from being frozen, we regulate the pressure of the gas phase of the volume of LN2 with a relief valve that vents at 2.5–3 bar ("R2" in figure 1). In stable conditions, the nitrogen volume is sealed by two electric valves, and the boiling temperature of the LN2 stays above 86 K.
@@ 2.2 Photon detector layout and readout
The layout of the SiPMs and the light source structure was designed with the primary goal of measuring the attenuation length and the yield of the LAr scintillation photon. The prototype of the mechanical structure for mounting the SiPMs and the photon source is shown in figure 3. The exterior diameter and the height of the cylindrical tower shown in the photo are adapted to the dimension of the OLAF LAr volume inner diameter and the attenuation length of the LAr scintillation photon in the order of some tens of centimeters. The tower has multiple rings each holding one SiPM, placed between 15 cm and 80 cm from the photon sources at the bottom layer of the tower. In figure 3, the photo on the left shows a prototype holder to host the SiPM (delivered by Hamamatsu with a ceramic housing) and to be fixed to a ring. The SiPMs face downward towards the photon sources, and their positions on the rings are staggered to avoid shadowing the SiPMs on the upper levels of the tower.
@@
To generate liquid argon scintillation photons of known intensity, we will use a fully encapsulated Am-241 source. The $\alpha$ particles from the primary alpha decay are shielded by the capsule, and only the accompanying 59.5 keV $\gamma$ (35.9% of the time) propagates in LAr with a mean free path of less than 1.5 cm. Using the coincidence of three SiPMs near the point where the $\gamma$ is expected to vanish, one can obtain a well-defined trigger for the position and time of the generated scintillation photons. The 3D rendering in the top-right of figure 3 illustrates the layout for the scintillation photon trigger. We will also place a green LED on the source holder for preliminary function validation and timing calibration of the SiPMs. The figure 3 bottom-right photo shows the LED glued with optical cement to the prototype holder, before the structure was completely submerged in LAr for testing.

### v2

@@ 1 Introduction
Liquid argon (LAr) has been used in neutrino and dark matter experiments as an active medium because of its excellent properties in charge yield and transport, as well as its capacity as a scintillator [1]. To detect LAr scintillation photons of the vacuum ultraviolet (VUV) wavelength range, standard method uses wavelength shifting materials to produce secondary photons in the visible wavelength range to which typical photomultipliers have the sensitivity.
@@
At Roma Tre university, we are preparing the Optical Liquid Argon Facility (OLAF), with a cylindrical volume of 40 L LAr to host VUV SiPMs and photon sources inside. The first of the two main goals of the facility is to characterize the optical properties of LAr as a scintillation detector to be used in the LEGEND-1000 experiment [6] as active shielding. Secondly, the facility is expected to function as a test bench for the R&D of the LAr scintillation detector in general. The compact size of the OLAF setup is important for a fast turnaround time to test different configurations and designs of the measurement. In the next section, we will introduce the experimental setup in detail.
@@ 2.1 Cryogenics for LAr
The central part of the cryogenic system is a cylindrical vessel made of stainless steel, about 2 m high, as shown in figure 2. The innermost vessel can host LAr up to 40 L, surrounded by a jacketed volume of about 20 L where the LN2 is filled. An external vacuum layer then thermally isolates the inner volumes from the surroundings. The three layers are welded together at the top part of the vessels, as shown in the photo in figure 2a. A 200 L Dewar is used to continuously provide LN2 to OLAF.
@@
The thermally insulated pipes were equipped, as shown in figure 2b, to vent the cold gas to the roof outside the laboratory. The top of the LAr volume is fitted with a transparent sight glass through which the interior can be visually inspected. Temperature sensors placed in multiple vertical locations indicate the liquid level of the LAr and LN2 volumes. The feed-through flanges as shown in figure 2a provide separate sockets for the SiPMs and the temperature sensors.
@@
When the LN2 volume of OLAF is filled, the thermal contact of the jacketed wall provides a cooling power of a few kilowatts to the argon volume. By controlling the flow of the argon gas, we can produce tens of liters of LAr in a few hours  [^3] This includes the time required to cool the vessels from room temperature with LN2. At 1 atm, the boiling temperature of LN2 (77 K) is lower than the freezing temperature of LAr (83 K). To prevent LAr from freezing, we regulate the pressure of the gas phase of the volume of LN2 with a relief valve that vents at 2.5–3 bar ("R2" in figure 1). In stable conditions, the nitrogen volume is sealed by two electric valves indicated by EV1 and EV2 in figure 1, and the boiling temperature of LN2 stays above 86 K.
@@ 2.2 Photon detector layout and readout
The layout of the SiPMs and the light source structure was designed with the primary goal of measuring the attenuation length and the yield of the LAr scintillation photon. The prototype of the mechanical structure for mounting the SiPMs and the photon source is shown in figure 3. The exterior diameter and the height of the cylindrical tower shown in the photo are adapted to the dimension of the OLAF LAr volume inner diameter and the attenuation length of the LAr scintillation photon of the order of some tens of centimeters. The tower has multiple rings each holding one SiPM, placed between 15 cm and 80 cm from the photon sources at the bottom layer of the tower. In figure 3, the photo on the left shows a prototype holder to host the SiPM (delivered by Hamamatsu with a ceramic housing) and to be fixed to a ring. The SiPMs face downward towards the photon sources, and their positions on the rings are staggered to avoid shadowing the SiPMs on the upper levels of the tower.
@@
To generate liquid argon scintillation photons of known intensity, we will use a fully encapsulated Am-241 source. The $\alpha$ particles from the primary alpha decay are shielded by the capsule, and only the accompanying 59.5 keV $\gamma$ (35.9% of the time) propagates in LAr with a mean free path of less than 1.5 cm. Using the coincidence of three SiPMs near the point where the $\gamma$ vanishes, one can obtain a well-defined trigger for the position and time of the generated scintillation photons. The 3D rendering in the top-right of figure 3 illustrates the layout for the scintillation photon trigger. We will also place a green LED on the source holder for preliminary function validation and timing calibration of the SiPMs. The figure 3 bottom-right photo shows the LED glued with optical cement to the prototype holder, before the structure was completely submerged in LAr for testing.